Isaac Lake 1790-1851 Elizabeth Bentall 1785-1825 m Isaac Lake 1811-1889  ?  Sarah Lake 1812-1900 Elizabeth Lake 1813-1901 Chamberlayne Lake 1815-1815  Chamberlayne Hickman Lake 1817-1905  William Bentall Lake 1819-?  Amelia Forsbrey  Mary Ann Lake 1820-1835  Emma Lake 1822-1877  Arthur Bentall Lake 1824-1889 Susannah Eleanor Coote 1824-1919 m  m  m Chamberlayne Hickman Lake  Martha ?  Emily ?  Joshua Lake m  m  Isaac Lake  Mary Caird Collyer ?-1923  Millie Lake  Joseph Forsbrey Lake  Elizabeth Coote m  m William Lake  Florence Stansbery  Anna ?  Fred Lake  John Read  Caroline Lake m  m  m Reg White  Doris Lake  ?  Margaret Lake m  m 3 children  Son  John Read  Cuthbert Lake ?-1947  Minnie Smith  Beatrice Lake  Harold Lake 1878-1939  Ralph Lake  Evelyn Lake  Hilda Lake m John Droven  Cynthia Lake  Anthony Chapman Veronica Lake  Robert Chambers  Audrey Lake m  m  m Catherine Droven  Herbert Droven  Nicholas Chapman Christopher Chapman  Roger Chapman  James Chambers  Frederick Lake 1851-1944  ?  Grace ?  ?  Nellie Middleton  Arthur Lake 1853-?  ?  Thomas Lamb  Eleanor Lake 1855-?  Emilie Lake 1856-1941  Florence Lake 1857-1934  George Lake 1859-1950  Laura Leppard  Ardley Lake 1860-1948  ? Smith  Hickman Lake 1862-?  Annie Davidson  Harry Spalding 1868-1910  Alice Lake 1863-1946  Osmond Lake 1865-?  Linda ? 1865-1949  Albert Lake 1867-1867  Ernest Coote Lake 1869-1953  Anna Maria Braund 1869-1929 m  m  m  m  m  m  m  m  m  m  m  m Daughter Anthony Lake  Marie ? m Peter Anthony Lake 1951-  Marion Lake  Richard Kastell  Muriel Lake m Richard Kastell  Brenda Rudgard  Michael Andrew  Robert Shackleton  Jean Kastell m  m  m Gregory Kastell 1955-  Simon Kastell 1957-  Arthur Lake  Humphrey Lake  ? Lee  Alice Lamb  Fred Lamb  ?  Gwen Lamb  Tom Godbold  Hilda Lamb  Nora Lamb  ? Allison  Blanche Lamb  Evelyn Lamb m  m  m 1921  m David Claud Lee  Anthony Lamb  ? m 1950  Osmond Lake ?-1918  George Lake  Joan ?  Irene Lake ?-1975 m  George Lake ?-1928  Blanche Norton ?-1945  Cyril Saunders  Elsie Lake ?-1946  Douglas Lake  Helen ? m  m  m Son  ?  Daughter  ?  Daughter Daughter Daughter m  m  2 children  Arthur Lake  ?  Kathleen Lake  Margery Lake  Mary Lake  Kenneth Lake ?-1969 m Hickman Lake  Kathleen Lake  Reginald Spalding 1894-1982  Yvonne Reid  Betty Gordon  Gerald Spalding 1895-1984  Alice Overall 1897-1985  Muriel Spalding 1898-1991 m  m  m David Phillips 1922-  Joan Spalding 1924-  Eric Bucknell 1922-  Lorina Spalding 1928- m  m David Phillips 1948-  Sharon Thomas 1956-  Graham Day 1949-  Patricia Phillips 1952- m  m Ian Phillips 1979-  Robert Phillips 1979-  Annette Day 1977-  William Day 1979-  Robert Bucknell 1949-  Claire Martin 1953-  John Bucknell 1951-  Jacqueline Harvey 1954- m  m Lydia Bucknell 1975-  Martin Bucknell 1977-  Edward Bucknell 1981-  Camilla Bucknell 1986-  Andrew Bucknell 1979-  Louise Bucknell 1980-  Christopher Bucknell 1984- Jock Lake  ?  Pat ?  Margery Lake m  m Son 1940-  Charles Ernest Lake 1905-1987 Marguerite Anna Coltart 1907-1960  Norma Jean Heichen 1922-  Lyall Theodore Oppen 1911-1997 Josephine Marjorie Lake 1908-1996 Arthur John Lake 1915-1936  Jessie Dykes m 1927  m 1950  m 1943  m 1936 Ernest George Lake 1928-  Mary Pauline Michajluk 1931-  James Robert Helps 1926- Marguerite Valerie Lake 1930- m 1952  m 1951  Kevin Anderson 1953-  Richard Frederick Matheson 1950-  Hilary Anne Lake 1952-  Douglas Charles Lake 1953-1986  Sheila Lee Mullen 1955-1995  John Bradford Lake 1958-  Sherrille Dean Paterson 1957-  Christine Margaret Ober 1966- m 1973  m 1980  m 1974  m 1986  m 1993 Randall George Lake 1954-  Robert Murray Lake 1957-  Ronald John Willis 1960-  Debra Marilyn Lake 1959- m 1985 Shannon Marie Willis 1988-  Alanna Michelle Willis 1989-  Kenneth James Helps 1952-  Joanne Catherine Carey 1955-  Helen ?? ?? ??-  James Thompson Greenough 1956-  Dianne Louise Helps 1955-  Brian Donald Alward 1959-  Laura Maureen Helps 1959- m 1974 m ????  m 1982  m 1979 Gavin James Helps 1978-  Shayna Alison Helps 1980-  Christopher James Greenough 1985-  Geoffrey Frederick Greenough 1986-  Amanda Valerie Greenough 1987- Leigh-Anne Marie Alward 1983-  Sean Michael Alward 1984-  Michael Thomas Alward 1988- Mark Samuel Matheson 1982-  Bradford David Matheson 1987-  Colleen Elizabeth Lake 1976-  Robyn Dawn Lake 1979-  Mitchell Bradford Lake 1998-  Arthur John Lake 1936-1953

1